Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework#1, to be discussed on Thursday, 10 January and Tuesday, 15 January. §4.5 #27, 29, 39, 41, 43, 45, 49, 51, 53, 59, 63, 67. §6.1 #7, 9 , 11, 19, 21, 23, 25, 37, 38.

• Homework A. Due Wednesday, 16 January 2002. Applications plus, page 341, #7. In addition answer the following question: d) Suppose that the Buffon needle experiment is carried out with a needle of length 10 cm and the spacing between the lines is 15 cm. We drop the needle 100 times and it lands on a line 43 times. Use this information to make an estimate for the value of π.

January 9, 2002